Whch of the following expressions represent the expandecd form of (3x-4y)^2?
A. 9x^2 + 24xy - 16y^2
B. 9x^2 - 24xy = 16y^2
C. 12x^2 + 16y^2
D. 12x^2 - 12xy = 16y^2

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

Whch of the following expressions represent the expandecd form of (3x-4y)^2?
A. 9x^2 + 24xy - 16y^2
B. 9x^2 - 24xy = 16y^2
C. 12x^2 + 16y^2
D. 12x^2 - 12xy = 16y^2
----------------
(3x-4y)^2 = 9x^2 - 24xy + 16y^2
That's none of the above.
A, B & D have equal signs ???

Answer by stanbon(75887)

Whch of the following expressions represent the expandecd form of (3x-4y)^2?
---------------
(3x-4y)^2 = 9x^2 - 24xy + 16y^2
============================================
Cheers,
Stan H.